OpenStudy (anonymous):

Show that f(x) = log(x+√x^2 +1 ) is an odd function .

OpenStudy (unklerhaukus):

Show \(f(-x)=-f(x)\)

OpenStudy (anonymous):

Do you know about the Rationalizing of Denominator??

OpenStudy (hba):

just put -x in place of x

OpenStudy (anonymous):

Guys we have to solve it further it is not easy as you are thinking.. We have to prove by rationalization process..

OpenStudy (anonymous):

yeah i know ..

OpenStudy (anonymous):

so what is the final answer ? .. can any one explain it more clearly .. please ?

OpenStudy (anonymous):

Find -f(x) first : \(-f(x) = -\log(x + \sqrt{x^2 + 1})\) --------1

OpenStudy (anonymous):

Now find \(f(-x)\)..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[f(-x) = \log(-x + \sqrt{x^2 + 1})\] Now we have to prove they are equal..

OpenStudy (anonymous):

.. okay

OpenStudy (anonymous):

\[-f(x) = -\log(x + \sqrt{x^2 + 1}) = \log(x + \sqrt{x^2 + 1})^{-1} = \log(\frac{1}{x + \sqrt{x^2 + 1}})\] \[\implies \log(\frac{1}{x + \sqrt{x^2 + 1}} \times \frac{x - \sqrt{x^2 +1}}{x - \sqrt{x^2 +1}})\] \[\implies \log(\frac{x - \sqrt{x^2 + 1}}{x^2 - x^2 - 1}) \implies \log(-x + \sqrt{x^2 + 1}) \implies f(-x)\] So, \[f(-x) = -f(x)\]

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

the function is niether odd nor even if it is odd then it should fill the following criteria f(-x)=-f(x) but it does not fulfill this criteria so i would say it it is niether odd nor even

OpenStudy (anonymous):

thank you sami for the help